OpenStudy (anonymous):

what is the exponential form of 0.045 ?

OpenStudy (anonymous):

e^2*10^-3

OpenStudy (anonymous):

how ?

OpenStudy (cwrw238):

scientific notation = 4.5 * 10^-2

OpenStudy (anonymous):

\[.045=\frac{ 45 }{1000 }=\frac{ 4.5*10 }{ 1000 }=\frac{ 4.5 }{100}=\frac{ 4.5 }{10^{2} }=4.5*10^{-2}\]